OpenStudy (anonymous):

What is the vertex of this quadratic function? f(x) = 2(x – 3)^2 – 5

OpenStudy (anonymous):

i think you wrote that wrong...check again

OpenStudy (anonymous):

yes :) thankyou

OpenStudy (anonymous):

\[f(x)=2(x-3)^2-5\implies f(x)=2x^2-12x+13\] now use \[x={-b\over2a}\]and solve:\[x={12\over2(2)}\] so you should end up with the vertex as \[x=3\]

OpenStudy (saifoo.khan):

Wait what?!

OpenStudy (anonymous):

what?

OpenStudy (saifoo.khan):

f(x) = 2(x – 3)^2 – 5 This is already in vertex form. a(x-h)^2 + k where (h,k) is the turning point.

OpenStudy (saifoo.khan):

Therefore: vertex = (3, -5)

OpenStudy (anonymous):

^or that...that is fine also

OpenStudy (saifoo.khan):

Like i said..

OpenStudy (anonymous):

haan haan zyada free na ho -.-

OpenStudy (saifoo.khan):

Mujhay yummydummy nay thori na kata k main free honga. :P :D

OpenStudy (anonymous):

^iska kya matlab??? :P meray kaatnay se free kyun?? LOL

OpenStudy (saifoo.khan):

lol, acha hua samajh nahi aaya.

OpenStudy (anonymous):

nahi nahi samjhao naa...kehtay kya ho?? :P

OpenStudy (saifoo.khan):

http://nooooooooooooooo.com/

OpenStudy (anonymous):

LOLOLOL